OpenStudy (anonymous):

the speed of a racing boat changes from 0 ft/s to 263 ft/s i six seconds. what is the acceleration of the boat in ft/s ? (the s in the last sentence has a 2 on top)

jimthompson5910 (jim_thompson5910):

acceleration = (change in speed)/(change in time)

jimthompson5910 (jim_thompson5910):

that formula will help you get the answer

OpenStudy (anonymous):

@jim_thompson5910 is that the answer because i dont get it

jimthompson5910 (jim_thompson5910):

what is the change in speed?

OpenStudy (anonymous):

0 ft/s to 264 ft/s in six seconds

jimthompson5910 (jim_thompson5910):

focus on the speeds for now

jimthompson5910 (jim_thompson5910):

what is the difference in those speeds?

OpenStudy (anonymous):

264 is greater then o ft/s ?

OpenStudy (anonymous):

i dont know im not very good in math \

jimthompson5910 (jim_thompson5910):

the change in speed is simply 263 - 0 = 263

jimthompson5910 (jim_thompson5910):

Change in time is 6-0 = 6 seconds

OpenStudy (anonymous):

ooh that was easy

jimthompson5910 (jim_thompson5910):

acceleration = (change in speed)/(change in time) acceleration = (263)/(6) acceleration = 43.833 don't forget the units for acceleration the 43.833 is approximate

OpenStudy (anonymous):

are we done with this problem @jim_thompson5910 ooh and thanks for helping me